A rectangle has a perimeter 36cm. If its length is 11cm what is its width?
Answer by jim_thompson5910(35256)

P = 2*(L+W)


36 = 2*(11+W)


36/2 = 11+W


18 = 11+W


18-11 = W


7 = W


W = 7


The width is 7 cm.